OpenStudy (anonymous):

You have a standard number cube. What is the probability of rolling a number less than 3, and then rolling a prime number? A.) 1/3 B.) 1/2 C.) 1/36 D.) 1/6

OpenStudy (anonymous):

a

OpenStudy (mathstudent55):

Numbers less than 3: 2, 3 Prime numbers: 2, 3, 5 P(number less than 3) = 2/6 = 1/3 P(prime) = 3/6 = 1/2 1/3 * 1/2 = 1/6